OpenStudy (anonymous):

Pre-Calc Question! PLEASE HELP Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Seven divided by x squared. + 10

OpenStudy (anonymous):

There are many ways. One example is \[f(x)=x+10,g(x)=\frac{ 7 }{ x^{2} }\]

OpenStudy (anonymous):

Thank you!